### v1

# PicoPose: Progressive Pixel-to-Pixel Correspondence Learning for Novel Object Pose Estimation

Abstract
Novel object pose estimation from RGB images presents a significant challenge for zero-shot generalization, as it involves estimating the relative 6D transformation between an RGB observation and a CAD model of an object that was not seen during training. In this paper, we introduce PicoPose, a novel framework designed to tackle this task using a three-stage pixel-to-pixel correspondence learning process. Firstly, PicoPose matches features from the RGB observation with those from rendered object templates, identifying the best-matched template and establishing coarse correspondences. Secondly, PicoPose smooths the correspondences by globally regressing a 2D affine transformation, including in-plane rotation, scale, and 2D translation, from the coarse correspondence map. Thirdly, PicoPose applies the affine transformation to the feature map of the best-matched template and learns correspondence offsets within local regions to achieve fine-grained correspondences. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC. PicoPose achieves state-of-the-art performance on the seven core datasets of the BOP benchmark, demonstrating exceptional generalization to novel objects represented by CAD models or object reference images. Code and models are available at https://github.com/foollh/PicoPose.
Index Terms— Perception for Grasping and Manipulation, Deep Learning for Visual Perception, Machine Learning for Robot Control
I Introduction
Object poses are typically represented by six degrees of freedom (DoFs) parameters, including 3D rotation and translation, to define the transformation from a canonical object space to the camera space. Estimating object poses is highly sought after in real-world applications, such as robotic manipulation [1, 2] and augmented reality [3, 4], and is therefore extensively explored in research.
Early research [5, 6, 7, 8] primarily focused on pose estimation with the same object CAD models for both training and testing phases, but lacked flexibility for the objects unseen during training. Later studies [9, 10, 11, 12, 13] addressed unseen objects within known categories by defining a normalized object coordinate space, but they still struggled with novel categories. With the advancement of foundation models [14, 15, 16], more recent research [17, 18, 19] has increasingly focused on handling entirely new objects to achieve zero-shot 6D object pose estimation, presenting a significant challenge for generalization.
Figure 1: An overview of our PicoPose with a three-stage pixel-to-pixel correspondence learning process for novel object pose estimation from RGB images. By progressively refining the correspondences, PicoPose significantly improves the accuracy of object poses computed via PnP/RANSAC.
For the zero-shot task of novel object pose estimation, recent methods using RGB-D images have achieved remarkable performance through techniques such as template matching with pose updating [17, 19] or point registration for pose computation [18]. The success of these approaches is largely attributed to the essential geometric support provided by depth maps, which supply crucial features for matching and offer geometric priors that enhance object localization in 3D space. However, the high cost of depth sensors often limits their practicality in real-world applications, making methods based solely on RGB images a more appealing option. Despite this, RGB-only approaches remain underexplored and generally fail to achieve competitive performance. Representative methods like GigaPose [20] and FoundPose [21], which rely on establishing correspondences between observed scenes and rendered templates via simply feature matching, often suffer from noisy correspondences prone to outliers, leading to imprecise pose predictions.
To this end, we introduce a novel framework for progressive pixel-to-pixel correspondence learning, termed as PicoPose, to enable precise pose estimation of novel objects from RGB images. As illustrated in Fig. 1, PicoPose progressively refines the correspondences between RGB observations and templates across three stages, significantly enhancing the accuracy of object poses computed from correspondences.
The architecture of PicoPose is illustrated in Fig. 2. More specifically, given an RGB image of a cluttered scene and a CAD model of an object that was not seen during training, PicoPose begins by rendering object templates from various viewpoints of the CAD model. These templates are then used in conjunction with zero-shot segmentation techniques (e.g., CNOS [22]) to detect the target object within the RGB scene. PicoPose then uses a three-stage correspondence learning process to identify the best-matched template for the detected object and to learn fine-grained pixel-to-pixel correspondences between them. Since each pixel in the template corresponds to a 3D surface point on the CAD model, we establish pairs of 2D positions on the observation and the corresponding 3D points on the template, which are subsequently used to compute the 6D pose through PnP/RANSAC. For the process of correspondence learning, the three stages are described as follows:
  • Stage 1: Feature Matching for Coarse Correspondences. In this stage, PicoPose utilizes a visual transformer backbone to capture features for matching the RGB observation and the rendered templates, identifying the best-matched one and obtaining coarse correspondences between the template and the observation.
  • Stage 2: Global Transformation Estimation for Smooth Correspondences. In this stage, PicoPose represents the coarse correspondences as a correspondence map, from which a 2D affine transformation, including in-plane rotation, scale, and 2D translation, is regressed to smooth the coarse correspondences and filter outliers.
  • Stage 3: Local Refinement for Fine Correspondences. In this stage, PicoPose applies the affine transformation to the feature map of the best-matched template and employs several offset regression blocks to learn correspondence offsets within local regions, progressively achieving fine-grained correspondences.
To assess the effectiveness of PicoPose, we train the model on the synthetic datasets of ShapeNet-Objects [23] and Google-Scanned-Objects [24] provided by [17], and test it on the seven core datasets of the BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V. The quantitative results on these datasets outperform existing methods, including GigaPose [20] and FoundPose [21], by a significant margin, demonstrating the zero-shot capability of PicoPose. We also successfully apply PicoPose in scenarios where object reference images, rather than object CAD models, are used to represent novel objects. Ablation studies are provided to demonstrate the advantages of individual designs in PicoPose.
In this paper, our key contributions are: (1) the introduction of PicoPose, a novel framework that leverages progressive pixel-to-pixel correspondence learning for pose estimation of novel objects from RGB images; (2) the development of three meticulously designed stages of correspondence learning to improve pose accuracy within PicoPose; and (3) the achievement of state-of-the-art results on the seven core datasets of the BOP benchmark for the RGB-based task.
II Related Work
Figure 2: An illustration of our proposed PicoPose. Given $N$ rendered object templates $\{\mathcal{T}_{i}\}_{i=1}^{N}$ for a target object and a detected RGB crop $\mathcal{I}$ of the observed scene, PicoPose progressively refines the correspondences through three meticulously designed stages to achieve precise pose estimation. In Stage 1, a vision transformer backbone is employed to extract features for template matching and correspondence building, identifying the best-matched template $\mathcal{T}$ and obtaining coarse correspondences between $\mathcal{I}$ and $\mathcal{T}$. In Stage 2, PicoPose encodes the learned correspondences into a correspondence map $\mathcal{A}$, from which it regresses a 2D affine transformation $\mathcal{M}$, including in-plane rotation $\alpha$, scale $s$, and 2D translation $(t_{u},t_{v})$, to smooth the coarse correspondences and filter out outliers. In Stage 3, the transformation $\mathcal{M}$ is applied to the feature map of $\mathcal{T}$, followed by the use of $L$ offset regression blocks to predict coordinate offsets $\Delta\mathcal{P}$ and a certainty map $\mathcal{C}$ within local regions, progressively enhancing the correspondence accuracy. Finally, the object pose is computed using PnP/RANSAC based on these refined correspondences.
II-A Methods Based on Image Matching
To address the generalization challenge, some approaches [26, 27, 17, 20, 19, 28] simplify the task of novel object pose estimation by using an image matching strategy, which involves rendering object templates in various poses and then retrieving the best-matched template to determine the corresponding pose. This strategy is often followed by downstream refinements, as in MegaPose [17] and GenFlow [28]. In contrast, FoundationPose [19] first updates the poses of templates before selecting the best-matched one.
II-B Methods Based on Pixel/Point Matching
This group of methods estimate object poses by establishing correspondences, including 2D-3D correspondences for RGB inputs and 3D-3D correspondences for RGB-D inputs. For instance, OnePose[29] matches the pixel descriptors in object proposals with the point descriptors obtained from Structure from Motion (SfM) to construct the 2D-3D correspondences, and OnePose++[30] further proposes coarse to fine matching to obtain more accurate correspondences. SAM-6D [18] learns 3D-3D correspondences through a two-stage point matching process incorporating background tokens. FoundPose [21] leverages the generalization capabilities of foundation models to extract pixel features and establish 2D-3D correspondences based on feature similarity.
III Method
III-A Overview of PicoPose
The goal of novel object pose estimation from RGB images is to determine the 6D transformation between an RGB observation and a CAD model of an object unseen during training. To address this task, we introduce a novel framework for generalizable pixel-to-pixel correspondence learning, termed as PicoPose, which enables precise estimation of the 6D transformation based on the established correspondences through PnP/RANSAC.
The architecture of PicoPose is illustrated in Fig. 2. For a given RGB image of a clustered scene and an object CAD model, we begin by rendering object templates from various viewpoints of the CAD model and employing zero-shot segmentation [22] to identify and crop the region containing the target object from the RGB scene. We then resize both the detected crop and the object templates to a fixed size of $H\times W$, representing them as $\mathcal{I}$ and $\{\mathcal{T}_{i}\}_{i=1}^{N}$, respectively, where $N$ is the number of templates. PicoPose utilizes these inputs to search for the best-matched template, denoted as $\mathcal{T}$, and progressively learns the pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$ across three stages, as detailed in Sec. III-B. Each foreground pixel on $\mathcal{T}$ corresponds to a 3D surface point on the CAD model, enabling us to establish pairs of 2D positions on $\mathcal{I}$ and their corresponding 3D points on $\mathcal{T}$, which are utilized for computing the 6D object pose via PnP/RANSAC.
III-B Progressive Pixel-to-Pixel Correspondence Learning of PicoPose
Given the resized detected crop $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and the object templates $\{\mathcal{T}_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$ from various viewpoints of the CAD model, PicoPose employs the ViT-L backbone [31], pretranined by DINOv2 [15], to extract their patch features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{M\times D}$ and $\{\mathcal{F}_{\mathcal{T}_{i}}\in\mathbb{R}^{M\times D}\}_{i=1}^{N}$, respectively, where $D$ is the feature dimension and $M=HW/196$ is the number of patches. PicoPose then identifies the best-matched templates $\mathcal{T}$ and performs three learning stages to build pixel-to-pixel correspondences between $\mathcal{I}$ and $\mathcal{T}$.
Stage 1: Feature Matching for Coarse Correspondences
Feature similarities between patches of the RGB observation $\mathcal{I}$ and the object templates $\{\mathcal{T}_{i}\}_{i=1}^{N}$, particularly with the best-matched template to $\mathcal{I}$, could establish coarse correspondences.
To enable more effective correspondence establishment, we initially retrieve the best-matched template $\mathcal{T}$ by scoring the degree of similarity between each template $\mathcal{T}_{i}$ and the observation $\mathcal{I}$. For each template $\mathcal{T}_{i}$, we obtain this template matching score $c_{i}$ by averaging the maximum feature cosine similarities of the foreground patches in $\mathcal{I}$ (identified by a prior-step zero-shot segmentation) with the patches in $\mathcal{T}_{i}$, which could be formulated as follows:
$\vspace{-0.1cm}c_{i}=\frac{1}{M^{\prime}}\sum_{j\in\texttt{FG}(\mathcal{F}_{\mathcal{I}})}\max_{k=1,\dots,M}\frac{<\mathbf{f}_{\mathcal{I},j},\mathbf{f}_{\mathcal{T}_{i},k}>}{|\mathbf{f}_{\mathcal{I},j}|\cdot|\mathbf{f}_{\mathcal{T}_{i},k}|},\vspace{-0.1cm}$ (1)
where $\mathbf{f}_{\mathcal{I},j}\in\mathcal{F}_{\mathcal{I}}$ and $\mathbf{f}_{\mathcal{T}_{i},k}\in\mathcal{F}_{\mathcal{T}_{i}}$ are the $j^{th}$ patch features in $\mathcal{I}$ and the $k^{th}$ patch features in $\mathcal{T}_{i}$, respectively, with <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denoting an inner product. $\texttt{FG}(\mathcal{F}_{\mathcal{I}})$ represents the indices of foreground patches in $\mathcal{I}$, and $M^{\prime}$ is the count of foreground patches.
From the set $\{\mathcal{T}_{i}\}_{i=1}^{N}$, the template $\mathcal{T}$ with the highest template matching score is chosen as the best match for establishing pixel-to-pixel correspondences. Furthermore, the feature similarities between $\mathcal{I}$ and $\mathcal{T}$ provide coarse correspondences by determining the most similar patch in $\mathcal{T}$ for each patch in $\mathcal{I}$.
Stage 2: Global Transformation Estimation for Smooth Correspondences
In Stage 1, we exploit feature matching to obtain the coarse and sparse correspondences between $\mathcal{I}$ and $\mathcal{T}$, which, however, often exhibit cluttered distributions with noise and many outliers, as shown in Fig. 1. Therefore, the objective of this stage is to improve the smoothness of these correspondences and filter out the outliers.
To achieve the objective, we adopt a global approach to estimate the 2D affine transformation $\mathcal{M}$ between $\mathcal{I}$ and $\mathcal{T}$, which can be parameterized with 4 degrees of freedom (DoFs) [20] as follows:
$\mathcal{M}$ $=\begin{bmatrix}s\cos(\alpha)&-s\sin(\alpha)&t_{u}\\ s\sin(\alpha)&s\cos(\alpha)&t_{v}\\ \end{bmatrix}\>,$ (2)
where $\alpha$ denotes the in-plane rotation angle, $s$ denotes the relative scale between $\mathcal{I}$ and $\mathcal{T}$, and $(t_{u},t_{v})$ represents the 2D translation of the object centroid in these two images. Applying $\mathcal{M}$ to transform the template $\mathcal{T}$ facilitates pixel alignments with $\mathcal{I}$, thus enabling smooth correspondences. We also note that combining $\mathcal{M}$ with the viewpoint rotation of $\mathcal{T}$ can give the 6D object pose.
As highlighted in Fig. 3, the correspondence map $\mathcal{A}$ between $\mathcal{I}$ and $\mathcal{T}$ can effectively capture the variations in $\alpha$, $s$ and $(t_{u},t_{v})$, thereby encapsulating the essential patterns for learning the affine transformation $\mathcal{M}$ and reducing the associated learning difficulties. Therefore, instead of directly concatenating the patch features $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ for regressing $\mathcal{M}$, we propose a more effective approach by utilizing the coarse correspondences obtained in Stage 1, represented as the correspondence map $\mathcal{A}$ at this stage, to realize the target.
More specifically, we first normalize the feature vectors of $\mathcal{F}_{\mathcal{I}}$ and $\mathcal{F}_{\mathcal{T}}$ to $\overline{\mathcal{F}}_{\mathcal{I}}$ and ¯$\overline{\mathcal{F}}_{\mathcal{T}}$, respectively, and compute the correspondence map $\mathcal{A}$ as $\mathcal{A}=\overline{\mathcal{F}}_{\mathcal{I}}(\overline{\mathcal{F}}_{\mathcal{T}})^{T}\in\mathbb{R}^{M\times M}$, where $M=HW/196$. Subsequently, $\mathcal{A}$ is reshaped to the size of $(H/14)\times(W/14)\times M$ and passed through several stacked convolutions to reduce the spatial dimensions to a global pose vector. Finally, three parallel Multi-layer Perceptrons (MLPs) are applied to the pose vector to learn $(cos(\alpha),sin(\alpha))$, $s$ and $(t_{u},t_{v})$ of $\mathcal{M}$, respectively. Fig. 2 gives the illustration on this stage.
Remark: In GigaPose [20], the affine transformation $\mathcal{M}$ is estimated from each coarse correspondence using a ResNet [32], rather than the visual transformer used in the prior step of template matching, and RANSAC is then applied for selecting the optimal $\mathcal{M}$. In contrast, in our Stage 2, the single estimate of $\mathcal{M}$ can be directly predicted based on the similarities of features extracted by the preceding visual transformer, eliminating the need for RANSAC as a post-processing step and providing auxiliary supervision for training the backbone.
Figure 3: Visualization of correspondence maps between the feature of a point on the RGB observation (marked by a yellow star) and the features of templates with various affine transformations, including in-plane rotation, scaling, and 2D translation.
Stage 3: Local Refinement for Fine Correspondences
With the 2D affine transformation $\mathcal{M}$ predicted in Stage 2, we can align the feature map of the best-matched template $\mathcal{T}$ with the RGB observation $\mathcal{I}$, thereby achieving the smooth correspondences. For any position $(u,v)$ on $\mathcal{I}$, the corresponding position $(u^{\prime},v^{\prime})$ on $\mathcal{T}$ could be obtained using $\mathcal{M}$ as follows:
$\begin{bmatrix}u^{\prime}\\ v^{\prime}\end{bmatrix}=\mathcal{M}\begin{bmatrix}u\\ v\\ 1\end{bmatrix}.\vspace{-0.2cm}$ (3)
The deviation between $(u,v)$ and $(u^{\prime},v^{\prime})$ can be interpreted as the commonly known “optical flow” [33]. For all pixels in $\mathcal{I}$, we compute their corresponding positions in $\mathcal{T}$, denoted as $\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$, via Eq. (3) to represent the smooth correspondences. The affine transformation applied to the feature map of $\mathcal{T}$, including rotation, scaling, and translation, is then achieved by using $\mathcal{P}$ as indices to gather features from $\mathcal{T}$, producing a transformed feature map aligned with $\mathcal{I}$. At this stage, we further learn the offsets $\Delta\mathcal{P}\in\mathbb{R}^{H\times W\times 2}$ to update $\mathcal{P}$ to $\mathcal{P}+\Delta\mathcal{P}$, enabling fine-grained correspondence adjustments within local regions.
We realize local refinement of correspondences in a progressively learning manner. Specifically, we first apply Dense Prediction Transformer (DPT) [34] to our backbone, generating $L$ hierarchical feature maps $\{\mathcal{F}_{\mathcal{I}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ and $\{\mathcal{F}_{\mathcal{T}l}\in\mathbb{R}^{H_{l}\times W_{l}\times D_{l}}\}_{l=1}^{L}$ of $\mathcal{I}$ and $\mathcal{T}$, respectively, where $H_{l}\times W_{l}$ denotes the spatial size and $D_{l}$ is the number of channels for the $l^{th}$ feature map. We then use $L$ offset regression blocks to iteratively update $\mathcal{P}$.
For the $l^{th}$ offset regression block, the current $\mathcal{P}$ is resized to $H_{l}\times W_{l}\times 2$ and scaled by dividing each 2D position within it by $[H/H_{l}$, $W/W_{l}]$ to ensure spatial consistency. We denote this resized and scaled version as $\mathcal{P}_{l}$, which we use as indices to gather features from $\mathcal{F}_{\mathcal{T}l}$, resulting in the transformed feature map $\mathcal{F}_{\mathcal{T}l}^{\prime}$ to align with $\mathcal{F}_{\mathcal{I}l}$. Additionally, we introduce a third correlation feature map $\mathcal{F}_{\mathcal{C}l}$ via a Correlation Lookup module, introduced in RAFT [35], to explicitly provide correlation degrees and facilitate easier learning of offsets; more details on this module can be found in the RAFT paper [35]. We then concatenate $\mathcal{F}_{\mathcal{I}l}$, $\mathcal{F}_{\mathcal{T}l}^{\prime}$, and $\mathcal{F}_{\mathcal{C}l}$ to form the input for two sequences of stacked convolutions, which are used to regress offsets $\Delta\mathcal{P}_{l}\in\mathbb{R}^{H_{l}\times W_{l}\times 2}$ and certainty map $\mathcal{S}_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$. $\Delta\mathcal{P}_{l}$ is then interpolated to the size of $H\times W\times 2$, scaled by multiplying the 2D coordinates within it by $[H/H_{l},W/W_{l}]$, and added to $\mathcal{P}$ for updating. The certainty map $\Delta\mathcal{S}_{l}$ represents the confidence of the regressed offsets and is upsampled to produce $\mathcal{C}_{l}^{\prime}\in\mathbb{R}^{H\times W}$.
With the $L$ offset regression blocks, we have the fine-grained $\mathcal{P}$, with a certainty map $\frac{1}{L}\sum_{i=1}^{L}\mathcal{C}_{l}^{\prime}$. For each foreground pixel in $\mathcal{I}$, if its correspondence certainty exceeds 0.5, we use the position in $\mathcal{P}$ to find the corresponding pixel in $\mathcal{T}$, which is linked to a 3D surface point. Therefore, all the pixel-to-pixel correspondences generate the associated 2D-3D pairs to compute the final object pose via PnP/RANSAC.
III-C Training of PicoPose
We perform end-to-end training of the three-stage correspondence learning process in PicoPose by optimizing the following objective:
$\min\mathcal{L}=\mathcal{L}_{coarse}+\mathcal{L}_{smooth}+\mathcal{L}_{fine},$ (4)
where $\mathcal{L}_{coarse}$, $\mathcal{L}_{smooth}$, and $\mathcal{L}_{fine}$ are the loss terms associated with each of the three stages, respectively.
In Stage 1, we adopt the InfoNCE loss [36], as used in GigaPose [20], as the training objective $\mathcal{L}_{coarse}$ to learn feature matching between the RGB observation $\mathcal{I}$ and the best-matched template $\mathcal{T}$.
TABLE I: Quantitative results of different methods on the seven core datasets of BOP benchmark [25]. We report the mean Average Recall (AR) among VSD, MSSD and MSPD, as introduced in Sec. IV-A. ‘FM’ denotes featuremetric pose refinement [21].
Method ##\##Hypothesis BOP Dataset Mean
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V
w/o Iterative Refinement
MegaPose [17] - 22.9 17.7 25.8 15.2 10.8 25.1 28.1 20.8
GenFlow [28] - 25.0 21.5 30.0 16.8 15.4 28.3 27.7 23.5
GigaPose [20] 1 27.8 26.3 27.8 21.4 16.9 31.2 27.6 25.6
FoundPose [21] 1 39.5 39.6 56.7 28.3 26.2 58.5 49.7 42.6
PicoPose (Ours) 1 46.3 39.7 53.6 36.4 31.0 66.5 58.7 47.5
GigaPose [20] 5 29.6 26.4 30.0 22.3 17.5 34.1 27.8 26.8
FoundPose w/o FM [21] 5 39.6 33.8 46.7 23.9 20.4 50.8 45.2 37.2
FoundPose [21] 5 42.0 43.6 60.2 30.5 27.3 53.7 51.3 44.1
PicoPose (Ours) 5 49.2 41.3 58.4 37.8 32.7 67.6 57.6 49.2
w/ Refiner of MegaPose [17]
MegaPose [17] 1 49.9 47.7 65.3 36.7 31.5 65.4 60.1 50.9
GigaPose [20] 1 55.7 54.1 58.0 45.0 37.6 69.3 63.2 54.7
FoundPose w/o FM [21] 1 55.4 51.0 63.3 43.0 34.6 69.5 66.1 54.7
FoundPose [21] 1 55.7 51.0 63.3 43.3 35.7 69.7 66.1 55.0
PicoPose (Ours) 1 60.5 56.6 63.6 46.5 40.1 75.9 68.7 58.8
MegaPose [17] 5 56.0 50.7 68.4 41.4 33.8 70.4 62.1 54.7
GigaPose [20] 5 59.8 56.5 63.1 47.3 39.7 72.2 66.1 57.8
FoundPose w/o FM [21] 5 58.6 54.9 65.7 44.4 36.1 70.3 67.3 56.8
FoundPose [21] 5 61.0 57.0 69.4 47.9 40.7 72.3 69.0 59.6
PicoPose (Ours) 5 61.1 57.1 65.0 48.2 42.1 76.3 69.6 59.9
In Stage 2, we predict the 2D affine transformation $\mathcal{M}$, including in-plane rotation angle $\alpha$, scale $s$, and 2D translation $(t_{u},t_{v})$, to generate smooth correspondences between $\mathcal{I}$ and $\mathcal{T}$. Letting $\hat{\alpha}$, $\hat{s}$, and $(\hat{t}_{u},\hat{t}_{v})$ represent the respective ground truths of the predicted parameters, we define the training objective $\mathcal{L}_{smooth}$ for this stage as follows:
$\mathcal{L}_{smooth}=\mathcal{L}_{geo}(\alpha,\hat{\alpha})+|\ln(s)-\ln(\hat{s})|+|t_{u}-\hat{t}_{u}|+|t_{v}-\hat{t}_{v}|,$ (5)
where $\mathcal{L}_{geo}(\alpha,\hat{\alpha})$ is the geodesic distance between two angles $\alpha$ and $\hat{\alpha}$, defined as follows:
$\mathcal{L}_{geo}(\alpha,\hat{\alpha})=\text{acos}\bigl{(}\text{cos}(\alpha)\text{cos}(\hat{\alpha})+\text{sin}(\alpha)\text{sin}(\hat{\alpha})\bigr{)}\>.$ (6)
In Stage 3, we employ the $L_{1}$ distance and the binary cross entropy objective to guide the learning of coordinate offsets $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and certainty maps $\{\mathcal{C}_{l}\}_{l=1}^{L}$ across all $L$ offset regression blocks as follows:
$\mathcal{L}_{fine}=\sum_{l=1}^{L}\lambda||\hat{\mathcal{C}}_{l}\cdot(\Delta\mathcal{P}_{l}-\Delta\hat{\mathcal{P}}_{l})||+\mu\mathcal{L}_{bce}(\mathcal{C}_{l},\hat{\mathcal{C}}_{l}),$ (7)
where $\{\Delta\hat{\mathcal{P}}_{l}\}_{l=1}^{L}$ and $\{\hat{\mathcal{C}}_{l}\}_{l=1}^{L}$ represent the corresponding ground truths of $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and $\{\mathcal{C}_{l}\}_{l=1}^{L}$, while $\lambda$ and $\mu$ are the weights to balance the loss terms. $\mathcal{L}_{bce}$ denotes the binary cross entropy objective. For the supervision of $\Delta\mathcal{P}_{l}$, we use the $\hat{\mathcal{C}}_{l}$ to mask and exclude invalid correspondences from the training process.
IV Experiments
IV-A Experimental Setups
Datasets we train the model on the synthetic datasets of ShapeNet-Objects [23] and Google-Scanned-Objects [24] provided by [17], using a total of $2,000,000$ training images. Evaluation is conducted on the seven core datasets of the BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V.
Implementation Details Following GigaPose [20], we utilize CNOS [22] for zero-shot segmentation and employ $N=162$ templates per object for template matching. The input images are resized to $H\times W=224\times 224$. We use $L=3$ stacked offset regression blocks to refine local correspondence, with feature maps at spatial resolutions of $16\times 16$, $32\times 32$, and $64\times 64$. The hyperparameters $\lambda=0.1$ and $\mu=1.0$ are applied in Eq. (7). PicoPose is trained using the ADAMW optimizer for 400,000 iterations, starting with an initial learning rate of $1\times 10^{-5}$, following a cosine annealing schedule. The training batch size is set to 64, distributed across 8 NVIDIA 3090 GPUs.
Evaluation Metrics We employ the standard BOP evaluation protocol [25] to report the mean Average Recall (AR) w.r.t three error functions, i.e., Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD) and Maximum Symmetry-Aware Projection Distance (MSPD). We also report the End-Point-Error (EPE), a widely used metric in flow estimation [35], to assess the quality of the correspondences.
Figure 4: Qualitative results of different methods without iterative refinement on the seven core datasets of BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V, arranged from left to right.
IV-B Comparisons with Existing Methods
We evaluate our proposed PicoPose against existing methods on the seven core datasets of the BOP benchmark [25]. Inspired by GigaPose [20], we enhance robustness of PicoPose by using the top 5 templates in Stage 2 and Stage 3 to learn fine correspondences and selecting the poses that best match these correspondences. The quantitative results are shown in Table I, where PicoPose significantly outperforms other methods, highlighting its superior zero-shot capability for novel object pose estimation through progressive correspondence learning. For example, a single model of PicoPose using top 5 templates outperforms the single models of GigaPose [20] and FoundPose [21] by $22.4\%$ and $5.1\%$ AR, respectively. In Table I, we also report results with the iterative refinement proposed by MegaPose [17], where PicoPose consistently outperforms the others, whether using 1 or 5 pose hypotheses for refinement. The visualizations in Fig. 4 further validate the advantages of PicoPose.
Results with Single Object Reference Images Since obtaining perfect object CAD models is not always practical, object images are sometimes used as references for pose estimation. Following GigaPose [20], we evaluate PicoPose in the most extreme scenario, i.e., with only one single reference image, by employing Wonder3D [37] to reconstruct the object CAD model from this image for pose estimation. As shown in Table II, PicoPose can successfully handle this extreme setting, achieving results comparable to GigaPose.
Runtime Analysis We report the per-image processing time,including segmentation and pose estimation, of different methods without iterative refinement in Table III. For a fair comparison, both GigaPose [20] and our PicoPose are tested on the same servers; as shown in Table III, PicoPose achieves comparable speeds while delivering more impressive results, demonstrating its accuracy and efficiency. Notably, while FoundPose [21] employs more advanced servers, it still incurs significantly higher computational costs when matching its extensive template library (800 templates per object). In contrast, both PicoPose and GigaPose maintain efficient performance while using only 162 templates per object.
TABLE II: Quantitative results of single reference image on LM-O.
Method ##\##Hypothesis AR
GigaPose [20] 1 17.9
PicoPose (Ours) 1 20.4
GigaPose [20] 5 18.3
PicoPose (Ours) 5 22.0
TABLE III: Per-image runtime of segmentation and pose estimation on LM-O. “##\##Hyp” denotes the number of pose hypotheses.
Method ##\##Hyp Server Time (s)
GigaPose [20] 5 NVIDIA V100 0.640
FoundPose [21] 5 Tesla P100 3.360
GigaPose [20] 1 0.631
GigaPose [20] 5 GeForce 1.209
PicoPose (Ours) 1 RTX 3090 0.659
PicoPose (Ours) 5 1.562
IV-C Ablation Studies and Analyses
We conduct ablation studies on LM-O, T-LESS and YCB-V datasets to evaluate the efficacy of designs in PicoPose. Except for specific cases, the results are achieved using only the best-matched templates in Stage 2 and Stage 3.
TABLE IV: Quantitative comparisons among different stages of correspondence learning.
Stage LM-O T-LESS YCB-V MEAN
6D Pose Estimation (AR $\uparrow$)
1 28.6 27.3 41.2 32.4
2 31.2 25.6 45.5 34.1
3 46.3 39.7 58.7 48.2
Translation Estimation (Accuracy $\uparrow$)
1 40.3 43.1 60.2 47.9
2 43.2 48.4 69.6 53.7
3 62.6 56.0 78.7 65.8
Correspondence Estimation (EPE $\downarrow$)
1 3.6 4.4 4.5 4.2
2 3.0 4.2 2.1 3.1
3 2.1 3.8 1.2 2.4
Efficacy of Progressive Correspondence Learning The key to the success of PicoPose lies in its design of progressive pixel-to-pixel correspondence learning. To evaluate this, we first analyze the quality improvements of correspondences and examine their impact on pose estimation. For Stage 2, the predicted affine transformations are applied to obtain correspondences. As shown in Table IV, pose precision improves with finer correspondences, supporting the core claim of this paper. Stage 2 smooths coarse correspondences and filters outliers (Fig. 5), significantly improving translation accuracy (errors $<5$ cm), while its overall 6D pose enhancement over Stage 1 is marginal, as Stage 2 does not include viewpoint rotation updates like Stage 1 with PnP. Stage 3 further enhances precision with a $14.1\%$ AR improvement by locally refining the correspondences from Stage 2.
TABLE V: Quantitative comparisons among different variants of Stage 2.
Method Pose Estimation (AR $\uparrow$) Processing Time on LM-O (s $\downarrow$)
LM-O T-LESS YCB-V MEAN Model Forward Post-Processing ALL
$\mathbf{F}_{ist}$ in GigaPose [20] 26.0 22.5 25.3 24.9 0.465 0.166 0.631
Stage 2 w/ concatenated features 27.0 21.7 35.0 27.9 0.329 - 0.329
Stage 2 w/ correspondence map 31.2 25.6 45.5 34.1 0.329 - 0.329
TABLE VI: Quantitative results with different initial correspondences inputted to Stage 3.
Initial Correspondence LM-O T-LESS
AR Time (s) AR Time (s)
Matching in Stage 1 34.5 0.655 25.0 0.612
Pose from Stage 1 45.7 0.936 33.9 0.823
Pose from Stage 2 46.3 0.659 39.7 0.617
Figure 5: Visualization comparisons between the coarse correspondences from Stage 1 and the smooth ones from Stage 2.
Efficacy of Stage 2 We first assess the effectiveness of Stage 2 by presenting the results without it in Table VI, where we initialize the input correspondences for Stage 3 in two ways: 1) by directly using the coarse correspondences from Stage 1, and 2) by using the predicted poses from Stage 1 via PnP/RANSAC to obtain smoother correspondences. The first approach yields less precise results due to the high noise in the correspondences, while the second is less efficient because of the additional PnP/RANSAC processing. Next, we conduct experimental comparisons with regression based on direct feature concatenation of the observation and the template. As shown in Table V, learning from correspondence maps proves to be more effective, as it explicitly models coarse correspondences and effectively captures the variations in affine transformations. Additionally, we replace our Stage 2 with $\mathbf{F}_{ist}$ from GigaPose [20], which is less efficient, as discussed in Sec. III-B.
Efficacy of Stacked Offset Regression Blocks in Stage 3 In Stage 3, we use $L$ offset regression blocks, specifically $L=3$ in our experiments with feature spatial sizes from DPT [34] set to $16\times 16$, $32\times 32$, and $64\times 64$, to refine correspondences within local regions. The results from different offset regression blocks are reported in Table VII. As shown in the table, performance progressively improves as more blocks are used, indicating that finer correspondences are achieved through progressive local refinement.
Influence of the Number of Templates We follow the setup of GigaPose [20] by using $N=162$ templates per object in our evaluation experiments. In Table VIII, we present additional quantitative results with different numbers of templates for both GigaPose and our proposed PicoPose. The results show improvement as more templates are used, since both methods rely on template matching to select the best-matched template for the target object. However, the rate of improvement slows as the number of templates increases. Notably, PicoPose is more effective than GigaPose when using fewer templates, further highlighting the advantages of PicoPose.
TABLE VII: Quantitative results of different offset regression blocks (denoted as “OR Block”) in Stage 3. “Size” denotes the spatial sizes of aligned feature maps for regression. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
OR Block Size LM-O T-LESS YCB-V MEAN
1 $16\times 16$ 33.4 27.2 42.3 34.3
2 $32\times 32$ 42.9 36.6 53.9 44.5
3 $64\times 64$ 46.3 39.7 58.7 48.2
TABLE VIII: Quantitative comparison with GigaPose [20] on the number of templates. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
Method ##\##Temp LM-O T-LESS YCB-V MEAN
GigaPose [20] 2 4.8 4.4 2.0 3.7
PicoPose 10.5 12.8 14.1 12.5
GigaPose [20] 6 11.5 9.2 5.9 8.9
PicoPose 27.5 25.0 39.5 30.7
GigaPose [20] 42 25.0 23.3 23.4 23.9
PicoPose 43.9 37.9 57.4 46.4
GigaPose [20] 162 29.6 26.4 27.8 27.9
PicoPose 46.3 39.7 58.7 48.2
V Conclusion
In this paper, we propose PicoPose, a novel framework for object pose estimation from RGB images that uses progressive pixel-to-pixel correspondence learning across three carefully designed stages. We demonstrate the zero-shot capabilities of PicoPose on seven core datasets of the BOP benchmark. In future work, we aim to further increase the speed of PicoPose to achieve real-time performance and explore ways to reduce its reliance on templates.